# vTrain: A Simulation Framework for Evaluating Cost-effective and Compute-optimal Large Language Model Training

### v1

@@ I Introduction
Transformer-based [66] large language models (LLMs) like ChatGPT [46], LLaMA [64, 65], and PaLM [9] are taking the world by storm. Driven by the rapidly increasing model parameter size as well as training token size, these large autoregressive AI models have demonstrated phenomenal performance on many tasks using a variety of evaluation protocols such as zero-shot, few-shot, and fine-tuning [4].
@@
In this work, we present a profiling-driven simulator which guides the evaluation of a cost-effective and compute-optimal LLM training system (aka Virtual Train, vTrain in short). As discussed above, exhaustively searching through the enormous design space of LLM’s training system hyperparameters over real AI training clusters is impractical given the enormous cost it will incur. Instead, we propose a profiling-based simulation methodology that accurately estimates the LLM training time of each design point in a high-performance manner, rendering the optimal LLM training system configuration to be determined in several tens of minutes using a high-end multi-core CPU server. While being extremely fast, the design of vTrain is driven by the following key observations that enable an accurate estimation of LLM’s training time. Latest AI algorithms are expressed as a direct acyclic graph where each graph node represents a neural network layer. In essence, the order in which the graph node gets executed determines the AI model’s execution dataflow. For LLM inference, it is challenging to accurately estimate the model’s execution dataflow because the autoregressive nature of LLMs render the LLM graph nodes’ execution order to be determined dynamically at runtime. In contrast, the execution order of LLM graph nodes for training is precisely defined at compile time so it is possible for vTrain to statically determine how many LLM graph nodes to execute and the order in which they are executed. Additionally, the execution time of each individual LLM graph node (each layer) over a target GPU architecture is highly deterministic and exhibits little variance across different runs. We employ the aforementioned properties of LLM training to develop a sophisticated profiling-based framework that translates how each individual LLM graph nodes (layers) are decomposed into low-level GPU CUDA kernels, in accordance to the LLM parallelization strategy employed, which we utilize to estimate the execution time of a given iteration of LLM training. To demonstrate vTrain’s practicality, we present several case studies tackling critical research problems of LLM training:
@@ II-B LLM Parallelization Strategies
GPipe [20] proposes pipeline parallel training which splits an input batch into smaller micro-batches, allowing the training iteration of different micro-batches to be pipelined. The waiting time for other workers (the pipeline bubbles) can be reduced by increasing the number of micro-batches but at the cost of increased memory footprint. To alleviate such overhead, PipeDream [33, 34] proposed an alternative schedule called one-forward-one-backward (1F1B), where the memory footprint of each worker is reduced by limiting the number of in-flight micro-batches to the pipeline depth (we revisit GPipe/1F1B later in Figure 7). In general, the latency incurred in pipeline parallelism’s inter-stage communication is negligible. This is because the Send-Receive communication simply exchanges data across two workers and is less sensitive to the interconnection bandwidth. Rather, overall performance is governed by the amount of pipeline bubbles and is thus more sensitive to the micro-batch size and the pipeline scheduling algorithm.
@@ III-A Simulation Framework Overview
Figure 7: 4-way pipeline parallelism using (a) GPipe and (b) 1F1B scheduling policy. Example assumes 4 micro-batches.
@@ III-C Profiling Module
The design objective of our profiling module is twofold: 1) determining the relationship between the high-level layer-node’s computation operator with the low-level CUDA kernels (tasks) and 2) profiling the execution time of each CUDA kernel over the target GPU device, which are utilized by vTrain’s simulation algorithm (Section III-E) to estimate LLM’s training time. The profiling module first executes all computation operators (e.g., forward pass of a MHA block) within the operator-granularity graph using the GPU, one-by-one, and collects the CUDA kernel traces (e.g., volta_sgemm_128x64_tn kernel for self-attention in MHA) at runtime using CUPTI (CUDA profiling tools interface) [43]. While CUPTI provides low-level information about the executed GPU kernels (e.g., names of the CUDA kernels and their execution time), it lacks the information regarding which operator each CUDA kernel is associated with. We employ the task-to-layer mapping methodology proposed by Zhu et al. [73] to identify all the low-level CUDA kernels associated with a particular high-level graph operator. Using this information, vTrain constructs an operator-to-task lookup table (Figure 4) which provides both the list of CUDA kernels associated with a given operator and the execution time of those kernels profiled over the target GPU.
A key challenge of profiling all the layer-node operators in the operator-granularity execution graph is its potentially high profiling overhead and slow simulation time. For an LLM with $L$ decoder layers trained with $N_{MB}$ micro-batches, the number of layer-nodes to profile and simulate becomes $O(L\times N_{MB})$. We make the key observation that practically all LLMs employ a model design methodology where an identically shaped decoder layer is stacked repeatedly with the same hyperparameters (h, L, s, n), each of which is partitioned evenly across all the GPUs for parallel training. This leads to each partitioned slice of the decoder layer, which is executed within a given GPU, to all consist of the same set of CUDA kernels, enabling us to drastically cut down on the set of operators to profile. Consider the operator-granularity execution graph in Figure 8. Because of LLM’s repetitive execution nature of identical decoder layer operators, it is sufficient to profile only a single Fwd MHA layer-node instead of profiling all Fwd MHA layer-nodes in every Fwd (i)s. In the rest of this paper, we refer to these operations as necessary operators (Figure 4). With such optimization in place, vTrain significantly reduces the number of operators to profile to $O(1)$. In Section III-F, we discuss vTrain’s profiling overhead as well as its simulation time for deployment.
@@
8:    for all $u\in V_{i}$ do
9:        if $u.ref=0$ then
10:           $Q\leftarrow Q\cup\{u\}$
11:        end if
12:    end for
13:end for
14:
15:while $Q\neq\emptyset$ do
16:    $u\leftarrow Q.pop()$triangleright Fetch a task in FIFO order
17:    $i\leftarrow u.device$
18:    $T[i]\leftarrow\max\left(T[i],u.start+u.duration\right)$triangleright Proceed the timeline
19:    for all $c\in u.child$ do
20:        $c.start\leftarrow\max\left(c.start,u.start+u.duration\right)$
21:        $c.ref\leftarrow c.ref-1$triangleright Update the child task
22:        if $c.ref=0$ then
23:           $Q\leftarrow Q\cup\{c\}$triangleright Update the task queue
24:        end if
25:    end for
26:end while
27:
28:$t_{iter}\leftarrow\max\left(T[i]\right)$
@@ III-E Simulation Algorithm for Estimating Training Time
With the low-level task-granularity execution graph in place, vTrain now has means to estimate an LLM’s single-iteration training time. Algorithm 1 provides a pseudo-code of how vTrain utilizes the task-level execution graph for estimating single-iteration training time. The simulation begins by initializing the timeline and the task queue. The timeline maintains the time advanced for each GPU within the training system. Before the simulation begins, our simulator first searches all tasks in the task-granularity graph that have no dependencies, and places those tasks into the task queue (line $4$-$11$). Afterwards, vTrain begins the simulation process to estimate training time, which goes through the following procedure: fetch a task from the task queue in FIFO order (line $14$), advance the corresponding timeline (line $16$), update the states of the child tasks (line $18$-$19$), and push newly executable tasks into the task queue (line $20$-$22$). The simulation is finished when there is no more tasks to execute at which point vTrain returns the predicted single-iteration training time. The total end-to-end wall clock training time to consume all training tokens is estimated by multiplying “predicted single-iteration training time” with “the total number of training iterations an LLM training needs to go through (i.e., total number of tokens /// tokens consumed per iteration)”.
@@ IV Simulation and Validation Methodology
Single-node training time validation. For intra-node single-iteration training time validation, we perform the actual training of LLMs using a single AWS EC2 p4d GPU instance which contains 8 NVIDIA A100 GPUs connected over NVLink/NVSwitch [40]. We collected 1,440 data points with various LLM model configurations and tensor/data/pipeline parallelization plans. NVIDIA’s NCCL [38] is used to profile and model key communication primitives of LLM training whose communication data size ranges from 1 MB to 1024 MB. The measured single-iteration training times of these data points over the 8 GPU system are compared against the vTrain predicted training time, showing a mean absolute percentage error (MAPE) of 19.66% with a coefficient of determination of 0.9587 (Figure 9(a)).
@@
Multi-node training time validation. Similar to single-node validation, we perform the actual training of various LLM model configurations and measure their single-iteration training times for our multi-node training validation. Our multi-node GPU system contains 64 multi-GPU server nodes where each node contains $8$ NVIDIA A100 GPUs, amounting to 512 GPUs. Intra-node communication is done over NVLink/NVSwitch while inter-node communication is conducted over four 200 Gbps Mellanox HDR InfiniBand HCAs. Due to the enormous monetary cost associated with LLM training at this scale, collecting sufficient amount of validation data for academic research in itself is an extremely challenging task – to the best of our knowledge, the only publicly available real-world measurements of large-scale LLM training (wall clock) time is reported in Microsoft-NVIDIA’s work on Megatron-Turing NLG [61] which only includes “3” LLM training data points. With the help of our collaborators in industry, we were able to secure $116$ data points for validating the aforementioned multi-node training time (i.e., total of $512$ A100 GPUs), which we compared against the vTrain-predicted training time. The validation data points were collected based on the LLM model configurations of Megatron-LM [36]. As discussed in Section III-D, we use a simple latency-bandwidth model to estimate inter-node communication time. Specifically, we calculate the communication latency of each All-Reduce operation (the most critical collective communication primitive in 3D parallel training, Figure 5) using the following analytical model based formula suggested in NVIDIA’s NCCL [41]:
@@
Here $t$ is the derived communication time, $S$ is the size of data being transferred, $n$ is the number of workers (i.e., GPUs) participating in the collective communication, and $B$ is the effective communication bandwidth. To better model the effective communication bandwidth, we introduce a tuning parameter called bandwidth effectiveness factor, denoted as $\alpha$, and express the effective communication bandwidth as $B=\alpha B_{max}$ where $B_{max}$ is the maximum interconnection bandwidth, i.e., 800 Gbps ($=$ 200 Gbps $\times$ 4). To identify the optimal $\alpha$ value, we sweep this value from 0.1 to 1.0 and compare the vTrain-predicted single-iteration training times vs. the measured single-iteration training time. We observed that the average prediction error was minimized at $\alpha=1.0$ in our evaluation setting, which assumes that all the inter-node network bandwidth can fully be utilized. Using an $\alpha$ value of 1.0, our multi-node training validation study showed a MAPE of 14.88% with a coefficient of determination of 0.9908 (Figure 9(b)), demonstrating that our simple latency-bandwidth model is sufficiently accurate for multi-node training time predictions.
@@ V-A Case Study #1: Cost-effective LLM Training Plan
We use the state-of-the-art $530$B Megatron-Turing NLG (MT-NLG) [61] model architecture and its parallelization strategy as a representative example to demonstrate vTrain’s feasibility in uncovering an alternative LLM training plan with better cost-efficiency. We focus on MT-NLG because it is most open to providing details on the hyperparameters of its training system configuration. Nonetheless, vTrain’s simulation methodology is equally applicable to other LLM training systems employing 3D parallelism. Under the baseline MT-NLG study, the authors propose a heuristic based 3D parallelism targeting the following LLM training scenario:
@@
[…]
In Figure 11, we re-plot the design space exploration results in Figure 10 by limiting the design points to those employing $8$-way tensor parallelism and summarize its single-iteration training time and the resulting GPU compute utilization. Since the baseline MT-NLG explores three design points with a total of 2,240/2,800/3,360 GPUs, we focus on those design points that require similar amount of GPU compute budget for a fair comparison. The results in Figure 11 show that vTrain is capable of uncovering several superior design points (the three red dots) than the baseline MT-NLG training systems (the three black dots), achieving higher cost-efficiency. We summarize such results in Table I where each row represents a given design point proposed in MT-NLG and the vTrain uncovered, more cost-effective LLM training system configurations. For instance, the last row in Table I compares the baseline MT-NLG vs. vTrain recommended designs, each requiring 3,360 GPUs vs. 2,880 GPUs, respectively. Notice how the design points uncovered with vTrain requires $14\%$ fewer GPUs, achieves $7.1\%$ higher GPU utilization, but only incurs $8.9\%$ longer end-to-end training time (i.e., additional two days). Using AWS EC2 P4d GPU instances [2] as a proxy to translate these numbers in monetary costs, our vTrain recommended design effectively reduces the training cost by $6.6\%$, from 10.08 to 9.41 million dollars, saving 0.67 million dollars.

### v2

@@ I Introduction
Transformer-based [73] large language models (LLMs) like ChatGPT [50], Gemini [70], LLaMA [72], and PaLM [9] are taking the world by storm. Driven by the rapidly increasing model parameter size as well as training token size, these large autoregressive AI models have demonstrated phenomenal performance on many tasks using a variety of evaluation protocols such as zero-shot, few-shot, and fine-tuning [4].
@@
In this work, we present a profiling-driven simulator which guides the evaluation of a cost-effective and compute-optimal LLM training system (aka Virtual Train, vTrain111vTrain is open-sourced at https://github.com/VIA-Research/vTrain. in short). As discussed above, exhaustively searching through the enormous design space of LLM’s training system hyperparameters over real AI training clusters is impractical given the enormous cost it will incur. Instead, we propose a profiling-based simulation methodology that accurately estimates the LLM training time of each design point in a high-performance manner, rendering the optimal LLM training system configuration to be determined in several tens of minutes using a high-end multi-core CPU server. While being extremely fast, the design of vTrain is driven by the following key observations that enable an accurate estimation of LLM’s training time. Latest AI algorithms are expressed as a direct acyclic graph where each graph node represents a neural network layer. In essence, the order in which the graph node gets executed determines the AI model’s execution dataflow. For LLM inference, it is challenging to accurately estimate the model’s execution dataflow because the autoregressive nature of LLMs render the LLM graph nodes’ execution order to be determined dynamically at runtime. In contrast, the execution order of LLM graph nodes for training is precisely defined at compile time so it is possible for vTrain to statically determine how many LLM graph nodes to execute and the order in which they are executed. Additionally, the execution time of each individual LLM graph node (each layer) over a target GPU architecture is highly deterministic and exhibits little variance across different runs. We employ the aforementioned properties of LLM training to develop a sophisticated profiling-based framework that translates how each individual LLM graph nodes (layers) are decomposed into low-level GPU CUDA kernels, in accordance to the LLM parallelization strategy employed, which we utilize to estimate the execution time of a given iteration of LLM training. To demonstrate vTrain’s practicality, we present several case studies tackling critical research problems of LLM training:
@@ II-B LLM Parallelization Strategies
GPipe [21] proposes pipeline parallel training which splits an input batch into smaller micro-batches, allowing the training iteration of different micro-batches to be pipelined. The waiting time for other workers (the pipeline bubbles) can be reduced by increasing the number of micro-batches but at the cost of increased memory footprint. To alleviate such overhead, PipeDream [37, 38] proposed an alternative schedule called one-forward-one-backward (1F1B), where the memory footprint of each worker is reduced by limiting the number of in-flight micro-batches to the pipeline depth (we revisit GPipe/1F1B later in Figure 7). In general, the latency incurred in pipeline parallelism’s inter-stage communication is negligible. This is because the Send-Receive communication simply exchanges data across two workers and is less sensitive to the interconnection bandwidth. Rather, overall performance is governed by the amount of pipeline bubbles and more sensitive to the micro-batch size and the pipeline scheduling algorithm.
@@ III-A Simulation Framework Overview
Figure 7: 2-way pipeline parallelism using (a) GPipe and (b) 1F1B scheduling policy. Example assumes 4 micro-batches.
@@ III-C Profiling Module
The design objective of our profiling module is twofold: 1) determining the relationship between the high-level layer-node’s computation operator with the low-level CUDA kernels (tasks) and 2) profiling the execution time of each CUDA kernel over the target GPU device, which are utilized by vTrain’s simulation algorithm (Section III-E) to estimate LLM’s training time. The profiling module first executes all computation operators (e.g., forward pass of a MHA block) within the operator-granularity graph using the target GPU, one-by-one, and collects the CUDA kernel traces (e.g., volta_sgemm_128x64_tn kernel for self-attention in MHA) at runtime using CUPTI (CUDA profiling tools interface) [47]. While CUPTI provides low-level information about the executed GPU kernels (e.g., names of the CUDA kernels and their execution time), it lacks the information regarding which operator each CUDA kernel is associated with. We employ the task-to-layer mapping methodology proposed by Zhu et al. [81] to identify all the low-level CUDA kernels associated with a particular high-level graph operator. Using this information, vTrain constructs an operator-to-task lookup table (Figure 4) which provides both the list of CUDA kernels associated with a given operator and the execution time of those kernels profiled over the target GPU. A key challenge of profiling all the layer-node operators in the operator-granularity execution graph is its potentially high profiling overhead and slow simulation time. For an LLM with $L$ decoder layers trained with $N_{MB}$ micro-batches, the number of layer-nodes to profile and simulate becomes $O(L\times N_{MB})$. We make the key observation that practically all LLMs employ a model design methodology where an identically shaped decoder layer is stacked repeatedly with the same hyperparameters (h, L, s, n), each of which is partitioned evenly across all the GPUs for parallel training. This leads to each partitioned slice of the decoder layer, which is executed within a given GPU, to all consist of the same set of CUDA kernels, enabling us to drastically cut down on the set of operators to profile. Consider the operator-granularity execution graph in Figure 8. Because of LLM’s repetitive execution nature of identical decoder layer operators, it is sufficient to profile only a single Fwd MHA layer-node instead of profiling all Fwd MHA layer-nodes in every Fwd (i)s. In the rest of this paper, we refer to these operations as necessary operators (Figure 4). With such optimization in place, vTrain significantly reduces the number of operators to profile to $O(1)$. In Section III-F, we discuss vTrain’s profiling overhead as well as its simulation time for deployment.
@@
8:    $Q\leftarrow Q\cup\{u|u\in V_{i},u.ref=0\}$
9:end for
10:
11:while $Q\neq\emptyset$ do
12:    $u\leftarrow Q.pop()$triangleright Fetch a task in FIFO order
13:    $i\leftarrow u.device$
14:    $T[i]\leftarrow\max\left(T[i],u.start+u.duration\right)$triangleright Proceed the timeline
15:    for all $c\in u.child$ do
16:        $c.start\leftarrow\max\left(c.start,u.start+u.duration\right)$
17:        $c.ref\leftarrow c.ref-1$triangleright Update the child task
18:        if $c.ref=0$ then
19:           $Q\leftarrow Q\cup\{c\}$triangleright Update the task queue
20:        end if
21:    end for
22:end while
23:
24:$t_{iter}\leftarrow\max\left(T[i]\right)$
@@ III-E Simulation Algorithm for Estimating Training Time
With the low-level task-granularity execution graph in place, vTrain now has means to estimate an LLM’s single-iteration training time. Algorithm 1 provides a pseudo-code of how vTrain utilizes the task-level execution graph for estimating single-iteration training time. The simulation begins by initializing the timeline and the task queue. The timeline maintains the time advanced for each GPU within the training system. Before the simulation begins, our simulator first searches all tasks in the task-granularity graph that have no dependencies, and places those tasks into the task queue (line $4$-$7$). Afterwards, vTrain begins the simulation process to estimate training time, which goes through the following procedure: fetch a task from the task queue in FIFO order (line $10$), then advance the corresponding timeline accordingly (line $12$). It is worth emphasizing that GPU kernel executions in LLM training are inherently sequential with few GPU kernels executed in parallel, due to the high computational intensity of these tasks. Despite such property, our simulation must account for the possibility of computation-communication overlap, as illustrated in Figure 5, which we faithfully model during the process in line $12$. Following this, the states of the child tasks are updated (line $14$-$15$), and newly executable tasks are pushed into the task queue (line $16$-$18$). The simulation is finished when there is no more tasks to execute at which point vTrain returns the predicted single-iteration training time. The total end-to-end wall clock training time to consume all training tokens is estimated by multiplying “predicted single-iteration training time” with “the total number of training iterations an LLM training needs to go through (i.e., total number of tokens /// tokens consumed per iteration)”.
@@ IV Simulation and Validation Methodology
Single-node training time validation. For intra-node single-iteration training time validation, we perform the actual training of LLMs using a single AWS EC2 p4d GPU instance which contains 8 NVIDIA A100 GPUs connected over NVLink/NVSwitch [44]. We collected 1,440 data points with various LLM model configurations and tensor/data/pipeline parallelization plans. NVIDIA’s NCCL [42] is used to profile and model key communication primitives of LLM training whose communication data size ranges from 1 MB to 1024 MB. The measured single-iteration training times of these data points over the 8 GPU system are compared against the vTrain predicted training time, showing a mean absolute percentage error (MAPE) of 8.37% with a coefficient of determination of 0.9896 (Figure 9(a)). One of the primary causes of vTrain’s single-node simulation error is rooted in how intra-node communication is modeled in our simulator. To model intra-node communication, vTrain profiles All-Reduce latencies while running NCCL primitives in an isolated environment. However, we observed that the latency of NCCL primitives during actual training is on average 30% higher than those measured in an isolated setting, a phenomenon also reported in [81]. Such discrepancy was especially more pronounced when tensor parallelism is employed. Specifically, vTrain underestimates the latencies of forward and backward passes of a tensor-parallel transformer layer because tensor parallelism exhibits more frequent All-Reduce operations (twice for each forward and backward passes of a transformer layer) than other parallelism strategies. Further improving vTrain to better accommodate such behavior for intra-node communication is left as future work.
@@
Multi-node training time validation. Similar to single-node validation, we perform the actual training of various LLM model configurations and measure their single-iteration training times for our multi-node training validation. Our multi-node GPU system contains 64 multi-GPU server nodes where each node contains $8$ NVIDIA A100 GPUs, amounting to 512 GPUs. Intra-node communication is done over NVLink/NVSwitch while inter-node communication is conducted over four 200 Gbps Mellanox HDR InfiniBand HCAs. The server nodes are connected in a two-level non-blocking fat-tree topology. Due to the enormous monetary cost associated with LLM training at this scale, collecting sufficient amount of validation data for academic research in itself is an extremely challenging task – to the best of our knowledge, the only publicly available real-world measurements of large-scale LLM training (wall clock) time is reported in Microsoft-NVIDIA’s work on Megatron-Turing NLG [67] which only includes “3” LLM training data points. With the help of our collaborators in industry, we were able to secure $116$ data points for validating the aforementioned multi-node training time (i.e., total of $512$ A100 GPUs), which we compared against the vTrain-predicted training time. The validation data points were collected based on the LLM model configurations of Megatron-LM [40]. As discussed in Section III-D, we use a simple latency-bandwidth model to estimate inter-node communication time. Specifically, we calculate the communication latency of each All-Reduce operation (the most critical collective communication primitive in 3D parallel training, Figure 5) using the following analytical model based formula suggested in NVIDIA’s NCCL [45]:
@@
Here $t$ is the derived communication time, $S$ is the size of data being transferred, $n$ is the number of workers (i.e., GPUs) participating in the collective communication, and $B$ is the effective communication bandwidth. To better model the effective communication bandwidth, we introduce a tuning parameter called bandwidth effectiveness factor, denoted as $\alpha$, and express the effective communication bandwidth as $B=\alpha B_{max}$ where $B_{max}$ is the maximum interconnection bandwidth, i.e., 800 Gbps ($=$ 200 Gbps $\times$ 4). To identify the optimal $\alpha$ value, we sweep this value from 0.1 to 1.0 and compare the vTrain-predicted single-iteration training times vs. the measured single-iteration training time. We observed that the average prediction error was minimized at $\alpha=1.0$ in our evaluation setting, which assumes that all the inter-node network bandwidth can fully be utilized. Using an $\alpha$ value of 1.0, our multi-node training validation study showed a MAPE of 14.73% with a coefficient of determination of 0.9887 (Figure 9(b)). The primary causes of vTrain’s multi-node simulation error are rooted on our simple latency-bandwidth model utilized for estimating inter-node communication latency, which introduces two main avenues of improvement. First, our inter-node communication model does not capture the effect of straggler GPU node’s training time at synchronization points, nor does it account for the latency overheads of NCCL kernel launches. Second, it is challenging for the latency-bandwidth communication model to accommodate the dynamic behaviors of a large, complicated network topology. For example, in Figure 3, there are four different data parallel groups across Node 0 and Node 3. Each group performs All-Reduce communications, and these communications from the four groups (depicted as gray arrows) share the same network building blocks (e.g., ToR switches), which can lead to potential interference between the groups. However, our analytical model is not able to capture such dynamic effects as it only accepts statically determined data size, number of GPUs involved in the inter-node communication, and link bandwidth as inputs. We believe the simulation errors of vTrain’s multi-node training time can be alleviated by incorporating the dynamic nature of inter-node communication into our analytical model or by employing existing sophisticated modeling techniques [75, 23]. We leave the augmentation of such advanced communication model as future work.
@@ V-A Case Study #1: Cost-effective LLM Training Plan
We use the state-of-the-art $530$B Megatron-Turing NLG (MT-NLG) [67] model architecture and its parallelization strategy as a representative example to demonstrate vTrain’s feasibility in rapidly identifying an LLM training plan with optimal cost-efficiency. We focus on MT-NLG because it is most open to providing details on the hyperparameters of its training system configuration. Nonetheless, vTrain’s simulation methodology is equally applicable to other LLM training systems employing 3D parallelism. Under the baseline MT-NLG study, the authors employ 3D parallelism targeting the following LLM training scenario:
@@
In order to train MT-NLG [67], the authors present three heuristic based LLM training plans they employed to balance training throughput and overall cost (the left half of Table I). Identifying a cost-effective LLM training plan requires countless trial-and-error exploration of various training plans in practice which, due to the enormous cost associated with such design space exploration, can only cover a subset of the full design space of 3D parallelism. In this section, we utilize vTrain to perform a full design space exploration of (t, d, p)-way 3D parallelism by sweeping through the design space up to tmax=16, dmax=32, and pmax=105-way tensor/data/pipeline parallelism. Figure 10 illustrates the result of such design space exploration in terms of MT-NLG model’s (a) single-iteration training time and (b) its GPU compute utilization. Not surprisingly, performance is best when the LLM training system employs a large number of GPUs, e.g., (16, 16, 105)-way 3D parallelism (Figure 10(a)). Unfortunately, such design point requires roughly 10×10$\times$10 × more GPUs than the baseline MT-NLG and also suffers from significant GPU compute underutilization (average $17\%$ GPU utilization, Figure 10(b)), worsening training cost-efficiency than all the baseline MT-NLG design points. Consequently, finding the most cost-optimal LLM training plan requires careful consideration of both training iteration time and GPU compute utilization.
@@
In Figure 11, we re-plot the design space exploration results in Figure 10 by limiting the design points to those employing $8$-way tensor parallelism and summarize its single-iteration training time and the resulting GPU compute utilization. Since the baseline MT-NLG explores three design points with a total of 2,240/2,800/3,360 GPUs, we focus on those design points that require similar amount of GPU compute budget for a fair comparison. The results in Figure 11 show that vTrain can effectively uncover LLM training plans (the three red dots) that are competitive to the baseline MT-NLG training systems (the three black dots). We summarize such results in Table I where each row represents a given design point (1) proposed in MT-NLG and (2) the vTrain uncovered, cost-effective LLM training system configurations (which was identified within several minutes). For instance, the first row in Table I compares the baseline MT-NLG vs. vTrain recommended designs, each requiring 2,240 GPUs vs. 2,016 GPUs, respectively. Notice how the design points uncovered with vTrain requires $10\%$ fewer GPUs, achieves $4.5\%$ higher GPU utilization, but only incurs $6.3\%$ longer end-to-end training time (i.e., additional two days). Using AWS EC2 P4d GPU instances [2] as a proxy to translate these numbers in monetary costs, our vTrain recommended design effectively reduces the training cost by $5\%$, from 9.01 to 8.62 million dollars, saving 0.39 million dollars. More critically, vTrain was able to identify such competitive LLM training plan within several minutes (i.e., exploring the full design space of Figure 10 takes less than 200 seconds), demonstrating the benefits our fast simulation framework provides to practitioners seeking to determine an efficient LLM parallelization and training plan across a distributed GPU system.
TABLE II: Validation of vTrain-predicted single-iteration training time vs. its actual single-iteration training time measured over our real $64$/$256$/$512$ A100 multi-GPU systems. $m$ refers to the micro-batch size. The design points highlighted in bold are LLM training plans newly uncovered by vTrain’s full design space exploration which achieve optimal cost-effectiveness (i.e., equivalent to the “Our findings” column in Table I), consistently providing lower single-iteration training time across all GPU sizes (the percentage of training time reduction is denoted inside the parentheses).
Param.
# GPUs $(t,d,p,m)$
Predicted
iteration time (sec)
Measured
iteration time (sec)
3.6 64 [40] (2, 32, 1, 16) 2.919 3.938
Ours (1, 64, 1, 8) 2.746 (6$\downarrow$) 3.567 (9$\downarrow$)
18.4 256 [40] (8, 32, 1, 4) 7.533 9.928
Ours (8, 32, 1, 8) 7.259 (4$\downarrow$) 9.604 (3$\downarrow$)
39.1 512 [40] (8, 32, 2, 4) 13.859 14.757
Ours (4, 32, 4, 2) 12.226 (12$\downarrow$) 13.876 (6$\downarrow$)
Because the LLM training costs reported in Table I are based on vTrain’s “simulated” single-iteration training time, we now quantitatively demonstrate how effective vTrain is in accurately estimating efficient training plans, cross-validating the vTrain simulated training time vs. the actual training time measured over real multi-node GPU systems. Validating the results of our design space exploration in Figure 10 and Table I over several thousands of GPUs is practically challenging for academic researchers due to the astronomical costs associated with such study. We therefore perform a design space exploration over a scaled-down version of MT-NLG that can be trained under 512 GPUs, the methodology of which we detailed below. In [40], the authors evaluate several scale-down versions of MT-NLG using 3D parallelism over distributed A100 GPU systems. These scale-down versions of MT-NLG maintain the identical, decoder-only transformer architecture but with reduced hyperparameters, such as smaller hidden size and reduced number of layers. Using these design points presented in [40], we take the following measure to validate how effective vTrain is in estimating efficient LLM training plans, the result of which is summarized in Table II. Table II shows two LLM training plans (1) those that are employed in [40] to train MT-NLG over $64$/$256$/$512$ A100 GPU systems, and (2) those that vTrain’s full design space exploration uncovered as providing better cost-effectiveness. The single iteration training time of these two LLM training plans (“[40]” and “Ours” in Table II) are then evaluated using both vTrain (“Predicted”) and real (“Measured”) multi-node GPU systems. As depicted in Table II, across all studied (64/256/512) GPU system configurations, the LLM training plan suggested by vTrain is able to consistently outperform those reported in [40], for “both” predicted and real training time measurements. These results highlight vTrain’s fidelity as well as its usefulness in cost-efficiently exploring the design space of LLM training plans in a fast and accurate manner.